2:10 DX adsec - idl zombie process during setFlat on DX before opening. Doug executed adsec stop/start on DX

2:43 Josh has pointed. Sending OSCO script to BD+70 622. Script is binocular with seeing limited on luci2 for collimation, and ACE Active on luci1 for OSCO. No data collection on luci2.
  • Seeing is 0.9-1.15 on the luci1 guide, and 1.5
  • Optimize gain failed with lots of safe skips, rotator 154, el 51. Safeskips failed optimize gain, during TT. made it through first pass just to fail
  • Second pass failed during TT optimization and adsec arb went down.
  • 2:56 resending preset after arb restarted.
  • OPtimize gain, lots of safe skips during ho1 optimization and put shell into failure
  • 2:03 adsec went into failure from the safeskips during optimize gain.
  • adsec stop start.
  • probem with tip tilt request for psf left side during set flat AO, set with 536 modes and Josh is clearing active optics on the left side.
  • 2:10 resending preset for OSCO. Last try...
  • Optimize gain fails, but without optimize gain the AUX loops does not updates and Apply Optical Gain fails. triicked it by starting Optimize gain and stopping through the WFS control gui. This tricked the FLAO gui into thinking Optimize gain was successful. Then tweaked gains manually, apply optical gain.
  • We ran 15 min over, so LBTB recorded this as downtime

-- JenniferPower - 15 Apr 2022
Topic revision: r2 - 15 Apr 2022, JenniferPower
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback